OpenStudy (anonymous):

simplify the expression, assume all variables are real numbers 5 sqaure root32(x+5)^5

hartnn (hartnn):

is it \(\huge \sqrt[5]{32(x+5)^5}\)

OpenStudy (anonymous):

i dont understand

hartnn (hartnn):

firsr write 32 as 2^5

OpenStudy (anonymous):

oh i see it now

OpenStudy (anonymous):

yes that is the correct format

hartnn (hartnn):

write 32 as 2^5 then use \(\sqrt[n]{ab}=\sqrt[n]{a}*\sqrt[n]{b}\)

OpenStudy (anonymous):

so would a be the 2^5?

hartnn (hartnn):

yes.

OpenStudy (anonymous):

and is n 5

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

yes, and b is (x+5)^5

OpenStudy (anonymous):

gotcha

hartnn (hartnn):

\(\large \sqrt[5]{32(x+5)^5}=\huge \sqrt[5]{2^5(x+5)^5}=\huge \sqrt[5]{...^5}*\huge \sqrt[5]{(...)^5}\) can u fill the blanks ...

OpenStudy (anonymous):

its 2 then the x+5 right

hartnn (hartnn):

yes.

hartnn (hartnn):

now, use , \(\huge \sqrt[n]{x^n}=x\)

OpenStudy (anonymous):

3125 is what i got...am i way off

hartnn (hartnn):

yes.

OpenStudy (anonymous):

ok what am i not doing?

hartnn (hartnn):

\(\large \sqrt[5]{32(x+5)^5}=\large \sqrt[5]{2^5}*\huge \sqrt[5]{(x+5)^5}=...(...)\) use, \(\huge\sqrt[n]{x^n}=x\)

OpenStudy (anonymous):

ok if i multiply 2 5 times then i get 3125

OpenStudy (anonymous):

then I would multiply that by the (x+5)^5?

hartnn (hartnn):

don't multiply 2 and 5, why do u multiply those ? what is \(\large \sqrt[5]{2^5}=???\)

OpenStudy (anonymous):

no, you know 2*2*2*2*2*=3125

OpenStudy (anonymous):

sorry, im horrible in math and I just dont get what im supposed to do next

OpenStudy (anonymous):

sorry i was multiplying 5's

hartnn (hartnn):

\(\huge\sqrt[n]{x^n}=x\) so, \(\large \sqrt[5]{2^5}=2\) did u get this ?

OpenStudy (anonymous):

not 2's

OpenStudy (anonymous):

how does it equal 2...

OpenStudy (anonymous):

if you dont mind explaining

hartnn (hartnn):

put x=2 , ....

OpenStudy (anonymous):

ok

hartnn (hartnn):

\(\large \sqrt[5]{(x+5)^5}=??\)

OpenStudy (anonymous):

so it would be 7+5 raised to the 5th power?

hartnn (hartnn):

how come 7 ?

OpenStudy (anonymous):

sorry 2, i threw us both off

hartnn (hartnn):

remenver this. \(\huge\sqrt[n]{x^n}=x\)

hartnn (hartnn):

so, \(\huge\sqrt[5]{2^5}=2\) and \(\huge\sqrt[5]{(x+5)^5}=x+5\) got this ??

OpenStudy (anonymous):

i do now..

hartnn (hartnn):

\(\large \sqrt[5]{32(x+5)^5}=2(x+5)=.....?\)

OpenStudy (anonymous):

if x is 2 that would mean that it would be 2(2+5) am I correct on that

hartnn (hartnn):

yes.

OpenStudy (anonymous):

so 14?

hartnn (hartnn):

yes. only if x=2

OpenStudy (anonymous):

lol ok...I'm confused again...I thought x does =2...no?

hartnn (hartnn):

is it given in question ?

OpenStudy (anonymous):

no x isnt given

hartnn (hartnn):

then its just 2(x+5) = 2x+10.

OpenStudy (anonymous):

ok I see it now, I understand so much better after i see the answer..I just never know how to get it...thanks for your help and your time

hartnn (hartnn):

practice more and more problems. welcome ^_^

OpenStudy (anonymous):

I'm going to, I was kind of practicing then, I just ran into one that my study plan didn't have examples for...well it did but it didn't give much details in the walk through

OpenStudy (anonymous):

@hartnn when I put the answer should I put the 2(x+5)=2x+10 or just 2x+10 because there is already an equal sign provided

hartnn (hartnn):

2x+10 is more simplified, so only 2x+10.

OpenStudy (anonymous):

alrighty thanks again